OpenStudy (anonymous):

f(x)=-6x^2+5x+3 A.-124,0,40 B.-6,5 C-0.83,-0.50 D.-2.00,1.67

OpenStudy (imstuck):

You are factoring, I presume? Do you know the quadratic formula?

OpenStudy (anonymous):

yes \[\frac{-b+-\sqrt{b^2-4(a)(c)} ? }{ 2(a) }\]

OpenStudy (imstuck):

So your a = -6, your b = 5, and your c = 3. Right? So fill them in.

OpenStudy (imstuck):

I did this three times and got the same factors all three times and none of them are your answers. I want to see what you get.

OpenStudy (anonymous):

Awe heck im suck at math thats why Im here to try and get help

OpenStudy (imstuck):

ok, but did you maybe copy something wrong? Either the equation or the answers? The first choice in your answers is a bit skewed. Check your typing and let me know. Ok? We can help you.

OpenStudy (imstuck):

If you can remember that quadratic formula, that's over half the battle here. The other half is getting the answers and equations correct. So check that for us.

OpenStudy (anonymous):

This work?

OpenStudy (imstuck):

Way better! Are you working on this one now?

OpenStudy (anonymous):

trying but I'm stuck math is extremely difficult for me I have a TBI

OpenStudy (imstuck):

I don't know what that is, but I can help you anyways, ok? Let's use that equation and fill in the quadratic formula with a = -5, b = 4, c = 9, ok?

OpenStudy (anonymous):

Traumatic Brain Injury ok

OpenStudy (imstuck):

\[\frac{ -4\pm \sqrt{(4)^{2}-4(-5)(9)} }{ 2(-5) }\]

OpenStudy (anonymous):

Im an Injure Marine

OpenStudy (anonymous):

injured*

OpenStudy (imstuck):

You'll get it; like I said you remembered the quadratic formula. That's a good thing! Sorry to hear you got hurt. : (

OpenStudy (anonymous):

8+180?

OpenStudy (imstuck):

\[\frac{ -4\pm \sqrt{16+180} }{-10 }\]

OpenStudy (anonymous):

188/-10?

OpenStudy (imstuck):

Be careful...4 squared is 16 not 8.

OpenStudy (imstuck):

\[\frac{ -4\pm \sqrt{196} }{ -10 }\]

OpenStudy (anonymous):

now divide that by 10?

OpenStudy (imstuck):

the square root of 196 has a perfect square in it...14. So that means we have this now: (it's a -10, don't forget)\[\frac{ -4\pm14 }{ -10 }\]

OpenStudy (imstuck):

There are two solutions in there now:\[\frac{ -4+14 }{ -10 }\]and\[\frac{ -4-14 }{-10 }\]

OpenStudy (imstuck):

That plus/minus sign is what gives you the two different--very different--solutions we get for the zeros.

OpenStudy (imstuck):

\[\frac{ -4+14 }{ -10 }=\frac{ 10 }{-10 }=-1\]

OpenStudy (imstuck):

\[\frac{ -4-14 }{ -10 }=\frac{ -18 }{ -10 }=\frac{ 9 }{ 5 }\]

OpenStudy (anonymous):

so how does the choices I gave you come into play?

OpenStudy (imstuck):

well, let's divide 9 by 5 and see what we get. 9/5 = 1.8. So we have x = -1.00 and x = 1.8

OpenStudy (imstuck):

See it there?

OpenStudy (anonymous):

right I got 1.8

OpenStudy (imstuck):

Cool! See! Not too bad...just keep your positives and negatives in order and you'll do a fine job!

OpenStudy (anonymous):

Im saying but that is no where on the choice you seen the file right?

OpenStudy (anonymous):

nvm sorry

OpenStudy (imstuck):

Right...it's the last choice on the second screenshot you sent. {-1.00, 1.80}

OpenStudy (anonymous):

you leave me :-(